OpenStudy (gamereternity):

what is the ratio x to y 3/5=x/x+2y

OpenStudy (gamereternity):

@John_ES

OpenStudy (gamereternity):

i need answers about to get timed out

rebeccaxhawaii (rebeccaxhawaii):

sorry we are not able to give direct answers :/

OpenStudy (john_es):

What is the correct problem? I find strange that x/x

OpenStudy (jdoe0001):

check http://openstudy.com/study/coc on tests and exams exams are meant to check if you studied, so, if you dunno, then you should have been asked this yesterday, not today

OpenStudy (gamereternity):

x over x+2y @John_ES

OpenStudy (john_es):

Ok you must solve the equation, \[\frac{3}{5}=\frac{x}{x+2y}\] but to find x/y

OpenStudy (john_es):

I give you a hint, \[3(x+2y)=5x\] Just using the cross multiplication. Do you know how to continue from this point?

OpenStudy (john_es):

You should just operate, \[3x+6y=5x\Rightarrow 6y=2x\Rightarrow 3=x/y\] I think this is what the problem is asking for.